OpenStudy (anonymous):

Hello, uh could you help me solve lim as u approaches 0 of (u^2cosu)/(sin2u)

OpenStudy (anonymous):

\[\lim_{u\to0}\frac{u^2\cos u}{\sin2u}~~?\]

OpenStudy (anonymous):

yes plz

OpenStudy (anonymous):

Recall the identity, \(\sin2u=2\sin u\cos u\). \[\begin{align*}\lim_{u\to0}\frac{u^2\cos u}{\sin2u}&=\lim_{u\to0}\frac{u^2\cos u}{2\sin u\cos u}\\\\ &=\frac{1}{2}\lim_{u\to0}\frac{u^2}{\sin u}\\\\ &=\frac{1}{2}\left(\lim_{u\to0}\frac{u}{\sin u}\right)\left(\lim_{u\to0}u\right)\end{align*}\]

OpenStudy (anonymous):

wait but what is the solution? is there no limit, also could you shoe with the limits approaching infinity and negative infinity please

geerky42 (geerky42):

\[\begin{align*}\lim_{u\to0}\frac{u^2\cos u}{\sin2u}&=\lim_{u\to0}\frac{u^2\cos u}{2\sin u\cos u}\\\\ &=\frac{1}{2}\lim_{u\to0}\frac{u^2}{\sin u}\\\\ &=\frac{1}{2}\left(\lim_{u\to0}\frac{u}{\sin u}\right)\underbrace{\left(\lim_{u\to0}u\right)}_{\Large \text{Here's the hint.}}\end{align*}\]

geerky42 (geerky42):

Put more effort in reading SithsAndGiggles's reply, please.

OpenStudy (anonymous):

thanks guys